Find the solution of value of x in the following1) x : -5/9=10/x

Grade:8

2 Answers

Arun
6 years ago
 
in  – 5/ 9 = 10 /x
if we cross multiply, we get
 – 5*x = 9*10
 – 5x = 90
x = – 90/5
x = –18        answer
hope this helps
ramya
5 years ago
we have -5/9=10/x
if we do cross multiplication then it will become -5x=90
then x=-90/5
so,x=-18 i hope this helps you